# On a novel UCP result and its application to inverse conductive scattering

### v3

@@ 1. Introduction
$\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}<\arg(x_{1}+\mathrm{i}x_{2})<\vartheta_{M}\},$ (1.1)
@@
$\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$
$\Gamma^{-}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{m},r\sin\vartheta_{m})^{\top},r>0\}.$
@@
Theorem 1.1 shows a novel type of UCP result associated with a PDEs system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
@@
Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\Omega\subset D$. Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given. Suppose that $\Omega\cap B_{r_{0}}={\mathcal{S}}_{r_{0}}$, where ${\mathcal{S}}_{r_{0}}$ is defined by (1.2), $r_{0}\in\mathbb{R}_{+}$ is sufficiently small and $\mathbf{0}\in\partial\Omega$. The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.2). Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDEs system:
@@
The PDE system described by (1.3) in Theorem 1.1 originates from the invisibility phenomena associated with the time-harmonic conductive medium scattering problem (1.7), which will be further elaborated upon in subsection 1.2. The direct scattering problem (1.7) involves solving for the scattered wave $u^{s}$ given the incident wave $u^{i}$ and the conductive medium scatterer $(\Omega,q,\eta)$, where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{2}$ with a connected complement such that $\Omega\Subset B_{R}\ (R\in\mathbb{R}_{+})$ and $u^{i}$ can be either a plane wave of the form
$u^{i}(\mathbf{x})=e^{{\rm i}k\mathbf{x}\cdot\mathbf{d}}$ (1.5)
@@
$u^{i}(\mathbf{x};\mathbf{z}_{0})=H_{0}^{1}(k|\mathbf{x}-\mathbf{z}_{0}|)$ (1.6)
@@
associated with a wave number $k\in\mathbb{R}_{+}$. Here, $\mathbf{d}\in\mathbb{S}^{1}$ is the incident direction of the plane wave, $H_{0}^{1}$ denotes the zeroth-order Hankel function of the first kind, and $\mathbf{z}_{0}$ signifies the location of the point source with $\mathbf{z}_{0}\in\mathbb{R}^{2}\setminus\overline{B}_{R}$ with $\Omega\Subset B_{R}$ and $R\in\mathbb{R}_{+}$. The function $q\in L^{\infty}(\Omega)$ characterizes the refractive index of the medium scatterer $\Omega$ with ${\rm supp}(q)\subset\Omega$, and the constant conductive parameter $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. On the other hand, the inverse problem corresponding to (1.7) aims to determine the shape of the underlying conductive medium scatterer $(\Omega,q,\eta)$ and its physical configuration $(q,\eta)$ through the measurement of the scattered wave. The novel UCP derived in Theorem 1.1 has several applications in the conductive medium scattering problem (1.7). Specifically, in Theorem 1.2, utilizing the UCP for (1.3), it can be shown that the conductive medium scatterer in (1.7) containing a convex irrational corner cannot be invisible. The detailed proof of Theorem 1.2 is postponed to Section 2.
Theorem 1.2.
@@
Consider the conductive medium scattering problem modeled by (1.7), associated with a given incident wave $u^{i}$ with the wave number $k$, as defined by (1.5) or (1.6). Let the bounded Lipschitz domain $\Omega\subset\mathbb{R}^{2}$ with a connected complement in (1.7) describe the corresponding conductive medium scatterer $(\Omega,q,\eta)$, where ${\rm supp}(q)\subset\Omega$, $q\in L^{\infty}(\Omega)$ and $\eta$ is a nonzero constant. If $\Omega$ has a convex irrational corner $\mathbf{x}_{0}$ with $q\in L^{\infty}(\Omega)\cap H^{2}(\Omega\cap B_{r_{0}}(\mathbf{x}_{0}))$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small, then $\Omega$ cannot be invisible for any wave number $k\in\mathbb{R}_{+}$.
@@
Using the unique continuation property (UCP) for (1.3) as established in Theorem 1.1, we can address the unique identifiability for the inverse problem of the conductive medium scattering problem (1.7). Specifically, in Theorem 1.3, we present both local and global unique determination of a conductive medium scatterer based on a single measurement of the scattered wave, where the single measurement refers to that taken under a fixed incident wave. The unique identifiability by a single measurement is challenging and has a long and colorful history in inverse scattering problems, often referred to as Schiffer problem in the literature; see [19, 18] and the references therein for more details. The detailed statement of Theorem 1.3 is provided in Theorem 2.1, along with the corresponding proof.
Theorem 1.3.
@@
Considering the conductive medium scattering problem (1.7), if the measurements of the two scattered wave corresponding to two conductive medium scatterers under a fixed incident wave are identical, then the difference between these two scatterers cannot contain a convex irrational corner. Furthermore, if two convex irrational polygonal conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, produce the same scattered wave measurements for a fixed incident wave, then $\Omega_{1}$ and $\Omega_{2}$ must be identical.
@@ 1.2. The conductive medium scattering and invisibility
Let $\Omega$ be a conductive medium scatterer suited in a homogenous isotropic background medium $\mathbb{R}^{2}$ with the medium configurations $q$ and $\eta$, which is a bounded Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$. The medium parameter $q\in L^{\infty}(\Omega)$ characterises the refractive index of the medium $\Omega$ with ${\rm supp}(q-1)\subset\Omega$ and the constant $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. In what follows, we write $(\Omega;q,\eta)$ to describe the conductive medium scatterer $\Omega$ with the physical configurations $q$ and $\eta$.
For a given incident wave $u^{i}$ impinging on $\Omega$, it generates a scattered wave $u^{s}$. The total wave $u$ is given by $u=u^{i}+u^{s}$. The incident wave $u^{i}$ can be either a plane wave of the form (1.5) or a point source of the form (1.6). The conductive medium scattering problem in $\mathbb{R}^{2}$ can be modeled as:
$\begin{cases}\Delta u^{-}+k^{2}qu^{-}=0&\text{in }\Omega,\\[5.0pt] \Delta u^{+}+k^{2}u^{+}=0&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] u^{+}=u^{-},\quad\partial_{\nu}u^{-}=\partial_{\nu}u^{+}+\eta u^{+}&\text{on }\partial\Omega,\\[5.0pt] u^{+}=u^{i}+u^{s}&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] \lim_{r\to\infty}r^{1/2}\left(\partial_{r}u^{s}-iku^{s}\right)=0,&\quad r=|\mathbf{x}|,\end{cases}$ (1.7)
@@
On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDEs system (1.11). If there exists a nontrivial pair of solutions $(w,v)$ to (1.11) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.11). Here, we would like to emphasise that throughout the article, $\eta$ can be taken to be identically zero, which reduces to the case with a normal scattering inhomogeneity of refractive index $q$, and all the results still hold equally. Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.11) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it.
In the following, we shall investigate the connection between the PDE system (1.3) in Theorem 1.1 and the conductive medium scattering (1.7) when the invisibility of the conductive medium scatterer $(\Omega;q,\eta)$ occurs. Recall that the incident wave $u^{i}$ is given by (1.5) or (1.6). Since $\Omega$ is bounded and has a connected complement, let $D:=B_{R}$ such that $\Omega\Subset B_{R}$, where $B_{R}$ is a disk centered at the origin with radius $R$. As discussed earlier, when $(\Omega;q,\eta)$ is invisible/transparent, we have $u^{\infty}\equiv 0$, and by Rellich’s theorem, it follows that $u^{s}\equiv 0$ in $D\backslash\overline{\Omega}$, which implies that $u^{+}\equiv u^{i}$ in $D\backslash\overline{\Omega}$. According to (1.7), we can readily derive:
@@ 1.3. Connections to existing results and main contributions
The UCP is a fundamental theory with various applications in solvability, controllability, optimal stability, and inverse problems for elliptic partial differential equations (cf. [18, 31, 32]). Several forms of UCP exist for second-order elliptic partial differential equations, including the weak UCP, strong UCP, UCP for Cauchy data, and UCP across a hypersurface [26, 50, 40, 4]. There are multiple approaches to establishing UCP for elliptic equations, such as doubling inequalities, Carleman inequalities, and three spheres inequalities. For further discussions on this topic, we refer to [16, 26, 25, 31, 32, 33, 35, 45, 48] and references therein for a historical overview. Recently, the UCP result for the (fractional) Laplacian equation in Euclidian space or a Riemannian manifold can be found in [27, 38]. For elliptic PDE systems, [46] proved that the zero set of least energy solutions for Lane-Emden systems with Neumann boundary conditions has zero measure. The weak UCP for a general second-order elliptic system was established in [30] under certain assumptions. Moreover, [49] derived UCP results for elliptic PDE systems, including strong UCP, weak UCP, and UCP for local Cauchy data.
@@
As discussed in the previous subsection, the invisibility or transparency in inverse scattering problems implies that the scatterer cannot be detected by external measurements, as the probing incident wave remains unperturbed and the corresponding scattered field is identically zero. However, when the scatterer exhibits geometrical singularities, such as corners or points of high curvature, as shown in [8, 12, 44, 24, 6] for time-harmonic acoustic scattering, invisibility or transparency cannot occur for any frequency associated with an incident plane wave. Recently, new methods based on free boundary problems have been introduced to investigate invisibility or transparency in isotropic and anisotropic acoustic medium scattering [13, 47, 14, 36].
@@
The quest to determine the geometrical and physical properties of scatterers through a single far-field measurement boasts a rich and extensive history in the realm of inverse scattering [2, 7, 18, 19, 34, 41]. This pursuit, often referred to as the Schiffer problem [19], aims to uniquely identify the location and shape of scatterers based on a single far-field measurement. When a scatterer is impenetrable, it is termed an obstacle; otherwise, it is considered a medium. In the inverse obstacle scattering, Lax and Phillips [37] demonstrated that the shape of a sound-soft obstacle can be uniquely identified through infinite far-field measurements. However, achieving uniqueness with just a single far-field measurement has thus far required additional a-priori knowledge of the geometry, such as the size or shape of the scatterer(cf. [20, 17, 41]).
@@
    We derive a novel type of UCP result for an elliptic second order PDEs system (1.3), where the components $v$ and $w$ of the solution $\mathbf{u}$ to (1.3) satisfy a transmission condition across two intersecting line segments within the underlying domain. These two intersecting line segments form an irrational corner.
@@
    We reveal that when the conductive medium scatter $(\Omega,q,\eta)$ has a convex irrational corner $\mathbf{x}_{0}$ and the physical parameter $q$ satisfies $H^{2}$ regularity near $\mathbf{x}_{0}$, then $\Omega$ radiates any wave number. Furthermore, we demonstrate that an admissible complex irrational polygonal scatterer can be uniquely determined using only a single far-field measurement, independent of its medium content.
@@
For the main contribution a), in contrast to the existing weak, strong, or Cauchy-type UCPs for elliptic systems in the literature, we consider the UCP for the case where the components of the solution to (1.3) fulfill a transmission boundary condition on two intersecting line segments within the domain, forming an irrational corner. This transmission boundary condition arises from the invisibility or transparency phenomena in conductive medium scattering, which holds both practical and physical significance. We will discuss this aspect in more detail in Section 2. To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDEs systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
For the main contribution b), we can substitute the technical condition on the a-priori information in [21], which requires the non-vanishing of the total wave at the corner points, with an irrationality condition related to the interior angles of the conductive polygonal medium $\Omega$. Although ensuring the non-vanishing nature of the total wave at the corner point is feasible in generic physical scenarios, the a-priori information on the irrationality of the interior angles of the underlying scatterer is less restrictive, making the new global unique identifiability more applicable in practice. We establish this uniqueness result by using the novel UCP result described in Theorem 1.1. Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ defined by (1.5) or (1.6).
@@
The remainder of the paper is organized as follows. In Section 2, we present the proofs of Theorems 1.1-1.3. In Section 3, we present our main results regarding the uniqueness of recovering the conductive polygonal-nest and polygonal-cell medium scatterers, including the scatterers shape, the partition of their interior polygonal-nest and polygonal-cell structure and the material parameters. Section 4 is devoted to the proofs of the main theorems of Section 3.
2. The proofs of Theorems 1.1-1.3
In this section, we aim to provide the proofs of Theorems 1.1-1.3. Specifically, we present a novel UCP result for an elliptic second-order PDEs system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement. To prove Theorem 1.1, we first introduce some auxiliary lemmas. Drawing upon the results in [12, 44], we present the following lemma, which constructs a complex geometrical optics (CGO) solution.
@@
[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2). If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
@@
By the interior regularity of the elliptic equation, we know that $v$ is analytic in $B_{r_{0}}$. When $v$ is analytic, it was proven in [22, Theorem 3.2] that $v(\mathbf{0})=0$. Using (2.6), one has

### v4

@@ 1. Introduction
$\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}\leq\arg(x_{1}+\mathrm{i}x_{2})\leq\vartheta_{M}\},$ (1.1)
@@
$\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$ (1.2a)
$\Gamma^{-}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{m},r\sin\vartheta_{m})^{\top},r>0\}.$ (1.2b)
@@
Theorem 1.1 shows a novel type of UCP result associated with a PDE system described by (1.4). The detailed proof of Theorem 1.1 is given in Section 2.
@@
Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\overline{\Omega}\subset D$. Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given. Suppose that $\Omega\cap B_{r_{0}}={\mathcal{S}}_{r_{0}}$, where ${\mathcal{S}}_{r_{0}}$ is defined by (1.3), $r_{0}\in\mathbb{R}_{+}$ is sufficiently small and $\mathbf{0}\in\partial\Omega$. The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.3). Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDE system:
@@
Remark 1.2.
The PDE system described by (1.4) in Theorem 1.1 originates from the invisibility associated with the time-harmonic conductive medium scattering problem (1.6). The phenomenon of invisibility in inverse scattering problems has been extensively studied (see, e.g., [7, 8, 14, 12, 13, 23, 34, 46, 53, 56] and the references therein). A direct result of Theorem 1.1 is Corollary 1.2, which establishes that if a medium scatterer contains a convex irrational corner, invisibility cannot occur. In other words, for a scatterer to be invisible while possessing a convex corner, the corner must necessarily be rational, not irrational. This conclusion aligns with the specific example provided in [46, Example 1.5], where invisible anisotropic medium scatterers feature a rational corner with an angle of the form $\ell\pi/m$ for integers $m\geq 2$ and $11\leq\ell<2m-$ 1. Thus, Corollary 1.2 is nearly optimal, as demonstrated by the existence of invisible scatterers with rational corners.
Moreover, the condition $\eta\neq 0$ in Theorem 1.1 can be interpreted as a nondegeneracy condition, which is necessary for establishing Theorem 1.1. Analogous nondegeneracy conditions have been proposed to characterize invisible medium scatterers; see [46, 12]. The nondegeneracy condition [46, Equation (1.6)] for the invisible medium scatterer requires a certain non-vanishing property associated with the incident wave in a neighborhood of a point on the boundary of the scatterer. This condition is used to prove that the boundary of the invisible medium scatterer is $C^{1,\alpha}$ around the underlying point. On the contrary, if the invisible medium scatterer has a corner, the nondegeneracy condition is not satisfied, as demonstrated in [46, Examples 1.5 and 1.7]. Such nondegeneracy conditions were also discussed in [12, Section 3.2]. Therefore, the nondegeneracy condition is necessary for characterizing invisible medium scatterers in [12, 46].
We then introduce the conductive medium scattering problem (1.6), which will be elaborated upon further in Subsection 1.2. Let $\Omega$ be a conductive medium scatterer suited in a homogenous isotropic background medium $\mathbb{R}^{2}$ with the medium configurations $q$ and $\eta$, which is a bounded Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$. The medium parameter $q\in L^{\infty}(\Omega)$ characterises the refractive index of the medium $\Omega$ with ${\rm supp}(q-1)\subset\Omega$ and the constant $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. In what follows, we write $(\Omega;q,\eta)$ to describe the conductive medium scatterer $\Omega$ with the physical configurations $q$ and $\eta$.
For a given incident wave $u^{i}$ impinging on $\Omega$, it generates a scattered wave $u^{s}$. The total wave $u$ is given by $u=u^{i}+u^{s}$. The incident wave $u^{i}$ is a solution to the Helmholtz equation. The conductive medium scattering problem in $\mathbb{R}^{2}$ can be modeled as:
[…]
The direct scattering problem (1.6) involves solving for the scattered wave $u^{s}$ given the incident wave $u^{i}$ and the conductive medium scatterer $(\Omega,q,\eta)$, where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{2}$ with a connected complement such that $\Omega\Subset B_{R}\ (R\in\mathbb{R}_{+})$ and $u^{i}$ can be a solution to the Helmholtz equation $\Delta u^{i}+k^{2}u^{i}=0$ at least in a neighborhood of $\Omega$, including, a plane wave of the form
$u^{i}(\mathbf{x})=e^{{\rm i}k\mathbf{x}\cdot\mathbf{d}}$
or a Herglotz wave with kernel $g_{j}\in L^{2}(\mathbb{S}^{1})$ of the form
$u^{i}(\mathbf{x})=\int_{\mathbb{S}^{1}}e^{\mathrm{i}k\xi\cdot\mathbf{x}}g_{j}(\xi)\mathrm{d}\sigma(\xi),\ \xi\in\mathbb{S}^{1},\ \mathbf{x}\in\mathbb{R}^{2}$
@@
$u^{i}(\mathbf{x};\mathbf{z}_{0})=H_{0}^{(1)}(k|\mathbf{x}-\mathbf{z}_{0}|)$
@@
associated with a wave number $k\in\mathbb{R}_{+}$. Here, $\mathbf{d}\in\mathbb{S}^{1}$ is the incident direction of the plane wave, $H_{0}^{(1)}$ denotes the zeroth-order Hankel function of the first kind, and $\mathbf{z}_{0}$ signifies the location of the point source with $\mathbf{z}_{0}\in\mathbb{R}^{2}\setminus\overline{B}_{R}$ with $\Omega\Subset B_{R}$ and $R\in\mathbb{R}_{+}$. The function $q\in L^{\infty}(\Omega)$ characterizes the refractive index of the medium scatterer $\Omega$ with ${\rm supp}(q)\subset\Omega$, and the constant conductive parameter $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. On the other hand, the inverse problem corresponding to (1.6) aims to determine the shape of the underlying conductive medium scatterer $(\Omega,q,\eta)$ and its physical configuration $(q,\eta)$ through the measurement of the scattered wave. The novel UCP derived in Theorem 1.1 has several applications in the conductive medium scattering problem (1.6). Specifically, in Corollary 1.2, utilizing the UCP for (1.4), it can be shown that the conductive medium scatterer in (1.6) containing a convex irrational corner cannot be invisible. The detailed proof of Corollary 1.2 is postponed to Section 2.
Corollary 1.2.
@@
Consider the conductive medium scattering problem modeled by (1.6), associated with a given incident wave $u^{i}$ satisfying the Helmholtz equation with wave number $k$. Let the bounded Lipschitz domain $\Omega\subset\mathbb{R}^{2}$ with a connected complement in (1.6) describe the corresponding conductive medium scatterer $(\Omega,q,\eta)$, where ${\rm supp}(q)\subset\Omega$, $q\in L^{\infty}(\Omega)$ and $\eta$ is a nonzero constant. If $\Omega$ has a convex irrational corner $\mathbf{x}_{0}$ with $q\in L^{\infty}(\Omega)\cap H^{2}(\Omega\cap B_{r_{0}}(\mathbf{x}_{0}))$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small, then $\Omega$ cannot be invisible for any wave number $k\in\mathbb{R}_{+}$.
@@
Using the unique continuation property (UCP) for (1.4) as established in Theorem 1.1, we can address the unique identifiability for the inverse problem of the conductive medium scattering problem (1.6). Specifically, in Corollary 1.3, we present both local and global unique determination of a conductive medium scatterer based on a single measurement of the scattered wave, where the single measurement refers to that taken under a fixed incident wave. The unique identifiability by a single measurement is challenging and has a long and colorful history in inverse scattering problems, often referred to as Schiffer problem in the literature; see [20, 19] and the references therein for more details. The detailed statement of Corollary 1.3 is provided in Theorem 2.1, along with the corresponding proof. Before that, we introduce the concepts of the irrational angle, irrational polygon and convex irrational polygon. A similar definition for irrational polygons can be founded in [16].
Definition 1.1.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a polygon. Let $\omega=\lambda\cdot\pi$, where $\lambda\in(0,2)$, be an interior angle of $\Omega$. The angle $\omega$ is termed irrational if $\lambda$ is an irrational number; otherwise, it is termed rational. Furthermore, if $\lambda\in(0,1)$ is an irrational number, the angle $\omega$ is referred to as a convex irrational angle. The polygon $\Omega$ is called an irrational polygon if all its interior angles are irrational; otherwise, it is considered rational. Additionally, the irrational polygon $\Omega$ is referred to as a convex irrational polygon if it is convex.
Corollary 1.3.
@@
Considering the conductive medium scattering problem (1.6), if the measurements of the two scattered wave corresponding to two conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, under a fixed incident wave are identical, then the difference between these two scatterers, $\Omega_{1}\aleph\Omega:=(\text{Co}(\Omega_{1})\setminus\text{Co}({\Omega_{2}}))\cup(\text{Co}(\Omega_{2})\setminus\text{Co}({\Omega_{1}}))$, cannot contain a convex irrational corner, where $\text{Co}(\Omega_{i}),i=1,2$ is the convex hull of $\Omega_{i}$. Furthermore, if two convex irrational polygonal conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, produce the same scattered wave measurements for a fixed incident wave, then $\Omega_{1}$ and $\Omega_{2}$ must be identical.
@@ 1.2. The conductive medium scattering and invisibility
On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDE system (1.10). If there exists a nontrivial pair of solutions $(w,v)$ to (1.10) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.10). Here, we would like to emphasise that throughout the article, $\eta$ can be taken to be identically zero, which reduces to the case with a normal scattering inhomogeneity of refractive index $q$, and all the results still hold equally. Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.10) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it. The boundary localization for an acoustic-elastic transmission eigenfunctions is revealed in [27].
In the following, we shall investigate the connection between the PDE system (1.4) in Theorem 1.1 and the conductive medium scattering (1.6) when the invisibility of the conductive medium scatterer $(\Omega;q,\eta)$ occurs. Recall that the incident wave $u^{i}$ satisfying the Helmholtz equation. Since $\Omega$ is bounded and has a connected complement, let $D:=B_{R}$ such that $\Omega\Subset B_{R}$, where $B_{R}$ is a disk centered at the origin with radius $R$. As discussed earlier, when $(\Omega;q,\eta)$ is invisible/transparent, we have $u^{\infty}\equiv 0$, and by Rellich’s theorem, it follows that $u^{s}\equiv 0$ in $D\backslash\overline{\Omega}$, which implies that $u^{+}\equiv u^{i}$ in $D\backslash\overline{\Omega}$. According to (1.6), we can readily derive:
@@ 1.3. Connections to existing results and main contributions
The UCP is a fundamental theory with various applications in solvability, controllability, optimal stability, and inverse problems for elliptic partial differential equations (cf. [19, 41, 42]). Several forms of UCP exist for second-order elliptic partial differential equations, including the weak UCP, strong UCP, UCP for Cauchy data, and UCP across a hypersurface [36, 58, 49, 5]. There are multiple approaches to establishing UCP for elliptic equations, such as doubling inequalities, Carleman inequalities, and three spheres inequalities. For further discussions on this topic, we refer to [17, 36, 35, 41, 42, 44, 45, 54, 57] and references therein for a historical overview. Recently, the UCP result for the (fractional) Laplacian equation in Euclidian space or a Riemannian manifold can be found in [37, 47]. For elliptic PDE systems, [55] proved that the zero set of least energy solutions for Lane-Emden systems with Neumann boundary conditions has zero measure. The weak UCP for a general second-order elliptic system was established in [40] under certain assumptions. Moreover, [52] derived UCP results for elliptic PDE systems, including strong UCP, weak UCP, and UCP for local Cauchy data. The UCP for the Maxwells system can be founded in [32].
@@
As discussed in the previous subsection, the invisibility or transparency in inverse scattering problems implies that the scatterer cannot be detected by external measurements, as the probing incident wave remains unperturbed and the corresponding scattered field is identically zero. However, when the scatterer exhibits geometrical singularities, such as corners or points of high curvature, as shown in [8, 14, 24, 53, 34, 7] for time-harmonic acoustic scattering, invisibility or transparency cannot occur for any frequency associated with an incident plane wave. Recently, new methods based on free boundary problems have been introduced to investigate invisibility or transparency in isotropic and anisotropic acoustic medium scattering [12, 56, 13, 46].
@@
The quest to determine the geometrical and physical properties of scatterers through a single far-field measurement boasts a rich and extensive history in the realm of inverse scattering [2, 6, 19, 20, 43, 50, 28]. This pursuit, often referred to as the Schiffer problem [20], aims to uniquely identify the location and shape of scatterers based on a single far-field measurement. When a scatterer is impenetrable, it is termed an obstacle; otherwise, it is considered a medium. In the inverse obstacle scattering, Lax and Phillips [48] demonstrated that the shape of a sound-soft obstacle can be uniquely identified through infinite far-field measurements. However, achieving uniqueness with just a single far-field measurement has thus far required additional a-priori knowledge of the geometry, such as the size or shape of the scatterer( cf.[21, 18, 50]). The unique identifiability for both inverse acoustic, elastic and electromagnetic impenetrable or penetrable scatterers under a-priori geometric knowledge using a few far-field measurements or single far-field measurement can be founded in [31, 25, 26, 33, 32]. Recently, the first stability estimating for simultaneous determination of the acoustic polygonal obstacle and the boundary impedance through a single far-field measurement is established in [30]. Furthermore, the effective medium theory for a time-harmonic electromagnetic inhomogeneous and possibly anisotropic medium with embedded obstacles can be founded in [29].
@@
    We derive a novel type of UCP result for an elliptic second order PDE system (1.4), where the components $v$ and $w$ of the solution $\mathbf{u}$ to (1.4) satisfy a transmission condition across two intersecting line segments within the underlying domain. These two intersecting line segments form an irrational corner.
@@
    We reveal that when the conductive medium scatter $(\Omega,q,\eta)$ has a convex irrational corner $\mathbf{x}_{0}$ and the physical parameter $q$ satisfies $H^{2}$ regularity near $\mathbf{x}_{0}$, then $\Omega$ radiates at any wave number. Furthermore, we demonstrate that an admissible complex irrational polygonal scatterer can be uniquely determined using only a single far-field measurement, independent of its medium content.
@@
For the main contribution a), in contrast to the existing weak, strong, or Cauchy-type UCPs for elliptic systems in the literature, we consider the UCP for the case where the components of the solution to (1.4) fulfill a transmission boundary condition on two intersecting line segments within the domain, forming an irrational corner. This transmission boundary condition arises from the invisibility or transparency phenomena in conductive medium scattering, which holds both practical and physical significance. We will discuss this aspect in more detail in Section 2. To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDE systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
For the main contribution b), we can substitute the technical condition on the a-priori information in [23], which requires the non-vanishing of the total wave at the corner points, with an irrationality condition related to the interior angles of the conductive polygonal medium $\Omega$. Although ensuring the non-vanishing nature of the total wave at the corner point is feasible in generic physical scenarios, the a-priori information on the irrationality of the interior angles of the underlying scatterer is less restrictive, making the new global unique identifiability more applicable in practice. We establish this uniqueness result by using the novel UCP result described in Theorem 1.1. Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ which is a solution to the Helmholtz equation with wave number $k$.
@@
The remainder of the paper is organized as follows. In Section 2, we present the proofs of Theorems 1.1 and Corollaries 1.2-1.3. In Section 3, we present our main results regarding the uniqueness of recovering the conductive polygonal-nest and polygonal-cell medium scatterers, including the scatterers shape, the partition of their interior polygonal-nest and polygonal-cell structure and the material parameters. Section 4 is devoted to the proofs of the main theorems of Section 3.
2. The proofs of Theorem 1.1 and Corollaries 1.2-1.3
In this section, we aim to provide the proofs of Theorems 1.1 and Corollaries 1.2-1.3. Specifically, we present a novel UCP result for an elliptic second-order PDE system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement. To prove Theorem 1.1, we first introduce some auxiliary lemmas. Drawing upon the results in [14, 53], we present the following lemma, which constructs a complex geometrical optics (CGO) solution.
@@
[19, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $B_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$. Then $v$ has the following spherical wave expansion in polar coordinates around the origin:
@@
By the interior regularity of the elliptic equation, we know that $v$ is analytic in $B_{r_{0}}$. When $v$ is analytic, it was proven in [24, Theorem 3.2] that $v(\mathbf{0})=0$. Using (2.6), one has